# Multi-objective Optimization in CPU Design Space Exploration: Attention is All You Need

### v1

@@ 5. Design of AttentionDSE
In this section, we introduce AttentionDSE, the first DSE framework that adopts the attention mechanism, suitable for high-dimensional DSE.
5.1. Overview of AttentionDSE
@@
Figure 3 provides an overview of AttentionDSE. It comprises three primary components: 1) Predictors, which provide rapid and accurate predictions of performance metrics (e.g., IPC) for given micro-architectures across benchmarks, thereby obviating the need for time-intensive simulations; 2) Micro-architecture Serialization, designed to reduce the computational burden associated with predictions, thereby expediting the performance evaluation process; 3) Bottleneck Analysis and Iterator, which identify the micro-architectural parameters that introduce performance bottlenecks within micro-architecture modules, facilitating the exploration of promising micro-architectural candidates by selectively modifying these parameters. The process of the AttentionDSE consists of a predictor training stage (①, ②, ③, ④) and an architecture exploration stage (❶, ❷, ❸, ❹, ❺).
Figure 3. Overview of the AttentionDSE.
Figure 4. The workflow of AttentionDSE framework in the exploration stage.
@@
The training stage is a one-time process, and the sequence of operations is represented by hollow numbers. This stage focuses on training predictors for performance metrics, ensuring that the models accurately capture the relationships between micro-architectural parameters and performance metrics. Unlike other approaches that rely on specialized sampling methods, AttentionDSE adopts a random sampling approach and then converts the design point to the sequence input, generating the training set (①, ②, ③). Next, the deep learning model, integrated with the PDA, functioning as the predictor, is trained to predict performance metrics (④). Benefiting from probabilistic attribute correlation in attention calculation, it’s more convenient to capture the intricate relationship between parameters, which informs exploration hints for the exploration stage.
@@
5.2. Attention-based Prediction Model
A fast and accurate performance predictor can replace traditional simulation methods, serving as the cornerstone of an efficient DSE framework. This approach significantly accelerates design iterations and reduces the overall design cycle. As discussed in Section 4, the attention-based model is a promising option for developing such a predictor. Neural networks possess strong nonlinear fitting capabilities, making them well-suited for learning from high-dimensional data. Additionally, they benefit from robust hardware acceleration and a well-developed GPU-based framework, collectively enhancing training and inference speed.
In this subsection, we propose an attention-based prediction model as the single-objective predictor. We aim to develop a more accurate and scalable predictor that emphasizes efficiency in high-dimensional DSE.
Input: $D$: Design space; $n$: The size of the training set; $k$: Embedding length; $Depth$: The depth of the model; $Epoch$: The total epochs for training.
@@ 5.2. Attention-based Prediction Model
3 Instantiate the prediction token $P_{t}$ with embedding length $k$;
@@ 5 Simulate $T$ to obtain $S_{ground\_truth}$ of IPC, Power, and Area;
6 for $epoch\leftarrow 0$ to $Epoch$ do
7       $S\leftarrow Serialize(T)$;
8       $S_{embed}\leftarrow Embedding(S)$;
9       $S_{embed}\leftarrow Concat(S_{embed},P_{t})$;
10       for $depth\leftarrow 0$ to $Depth$ do
@@ 12             $S_{mlp}\leftarrow MLP(S_{attn})$;
15      $S_{attn},A_{score}\leftarrow SelfAttention(S_{mlp})$;
16       Extract processed prediction token $S_{attn}\leftarrow S_{attn}[0,:]$;
17       $S_{pred}\leftarrow FullyConnectedLayer(S_{attn})$;
@@ 18       $BackwardPropagation(S_{pred},S_{ground\_truth})$;
The training procedure of the attention-based prediction model is demonstrated in Algorithm. 1. The prediction model consists of two key components: embedding preprocessing and the attention-based model. For embedding preprocessing, the procedure involves three steps: (a) Embedding, which transforms the discrete and heterogeneous alternatives of each micro-architectural parameter into a continuous numerical representation, simplifying optimization and enhancing the representation of relationships within the micro-architecture; (b) Position Embedding (PE), which assigns positional encoding to each item in the parameter sequence, enabling the model to learn parameter attributes more effectively and constrain the embedding space of each parameter; and (c) Prediction Token (PT), which gathers information from parameter embeddings using the self-attention mechanism and fuses it through multilayer perceptron (MLP), allowing the model to perceive the entire parameter sequence effectively. As for the attention-based model, to ensure the generalizability and versatility of the model, the predictor includes only vanilla self-attention, MLP, and normalization operations, which are the minimal necessary components of the transformer model (transformer, ). This streamlined design aims to maintain the attention mechanism’s inherent advantages. By focusing on these core components, we can leverage the self-attention mechanism’s ability to capture information within the input design point, allowing for the effective handling of complicated relationships between the design point and performance metrics in high-dimensional DSE tasks. The inclusion of MLP layers provides the model with the capacity to learn complex patterns and representations, while normalization operations ensure stable and efficient training by mitigating issues related to internal covariate shifts.
5.3. Perception-driven Attention Mechanism
@@
In previous work (GPR, ; MoDSE, ; ensemble_learning, ; BOOMExplorer, ), the order of parameters is not a concern. Because GPR (GPR, ; BOOMExplorer, ) relies on the distance and similarity between input data rather than their arrangement order in statistical regression methods. Similarly, in ensemble learning (ensemble_learning, ; MoDSE, ), the training process of multiple weak learners introduces randomness, making the order of parameters irrelevant. To address this challenge in the attention-based predictor, we introduce a serialization approach named Perception-driven Serialization (PDS), aimed at preserving architectural information and laying the groundwork for the sliding window attention technique. This technique employs a graph-based method to delineate relationships among all parameters within each pipeline stage.
Figure 5. The Perception-driven Serialization approach. (a) A Sample OoO CPU core; (b) Rename Stage; (c) Stage abstraction; (d) The steps of Perception-driven Serialization approach.
An example is illustrated in Fig. 5(a). We chose the Rename stage, shown in Fig. 5(b), as an example, where there are four parameters to be serialized including the free list (F), the rename map (R), the scoreboard (S), and the history buffer (H). The PDS first transfers the micro-architecture into a perceptual graph, which is an undirected graph based on the dataflow among these components, as depicted in Fig. 5(c). In a perceptual graph, vertices represent the parameters within each pipeline stage, while edges denote the datapaths between these components. The edges are categorized into two types: red edges, which connect to other pipeline stages and are labeled as external edges, and black edges, which connect to components within the current pipeline stage and are labeled as internal edges. To quantify the perception of each component, we leverage the perception degree $D$ of each parameter in the perceptual graph. The perception degree is calculated as follows:
(4) $D(\text{parameter})=\sum_{e\in E_{\text{parameter}}^{\text{internal}}}e-\sum_{e\in E_{\text{parameter}}^{\text{external}}}e$
@@
Figure 6. The attention weight calculation. (a) Vanilla self-attention mechanism; (b) PDA mechanism.
After the micro-architecture has been serialized, the vanilla self-attention mechanism can be replaced by the sliding window attention mechanism due to the parameter clusters. Fig. 6 shows an example of attention weight calculation compared with vanilla self-attention and PDA. For vanilla self-attention, every parameter must calculate the attention weight with every other parameter, as shown in Fig. 6(a). To optimize this process, we set the window size to the maximum perception degree among all parameters. In this example, the window size is set to five, as illustrated in Fig. 6(b).
In general, with the help of the PDA, each parameter calculates the attention weight only with its relative parameters. In this way, for a parameter vector of length $n$, the computational complexity is reduced from $O(n^{2})$ to $O(n)$. Thus, the PDA approach alleviates the limitations posed by number of parameters, enhances scalability, and reduces the training and inference time of the prediction model.
5.4. Attention-aware Bottleneck Analysis
Bottleneck analysis is a critical step in CPU design, enabling architects to identify the key factors that impact system performance. In this subsection, we propose the ABA algorithm, which equips AttentionDSE with the capability to analyze attention weight heatmaps generated by an attention-based predictor. The ABA algorithm, detailed in Algorithm 2, reveals the intricate relationships between micro-architectural parameters and performance metrics, thus guiding the exploration of design space. This approach ensures that the most influential design parameters are prioritized during optimization, leading to more efficient and effective DSE.
@@ 5       Set search queue $Q\leftarrow\varnothing$;
7             $x^{\prime}\leftarrow BottleneckAnalysis(x)$;
@@ 15      $\Omega\leftarrow GetParetoOptimal(Predictor(Q))$;
In the attention weight heatmap, each row represents how the current parameter is influenced by other parameters, reflecting the correlations between them. For a particular parameter, the sum of all weights in its corresponding column indicates the overall influence of this parameter on others. Therefore, if the total sum of attention weights in the column corresponding to a certain parameter is minimal for IPC or maximal for Power or Area, we infer that this parameter has a significant impact on the current performance metrics. ABA algorithm then updates the Pareto optimal set by updating the design point based on this bottleneck analysis, ensuring continuous improvement and exploration of the design space.
Figure 7. The example of ABA algorithm updating the micro-architectural parameters.
@@ 6. Experimental Setup
Simulation. We utilize GEM5 (gem5, ) as the timing-accurate simulator and McPAT (mcpat, ) as the power and area modeling tool. We extend GEM5 using Python to generate different CPU cores with varying configurations. To bolster the credibility of our experiments, we employ SPEC CPU2017 (SPEC2017, ) as the evaluation benchmark. For identifying resource utilization status, we utilize Simpoints (simpoints, ) for each workload evaluation. Each workload is divided into at most 30 clusters, with each cluster containing ten million instructions.
Baselines. Our baselines include the ActBoost (ActBoost, ), the BOOMExplorer (BOOMExplorer, ), and the MoDSE (MoDSE, ). Additionally, we implement ArchExplorer (ArchExplorer, ) for comparison with our ABA algorithm. The baselines represent recent SOTA work. All experiments are conducted on the Out-of-Order (OoO) CPU implemented by Gem5. The design space of the OoO processor is listed in Table 1. The values in the third column are formatted as “start number: end number: stride”. A two-level cache hierarchy with an 8192MB DRAM main memory configuration is employed. The size of the design space exceeds $6.89\times 10^{35}$, making exhaustive search impractical.
Table 1. Microarchitecture design space specification.
@@
[…]
@@
(6) $\mathrm{PHV}(\Omega)=\int_{\mathrm{X}}\mathbb{1}[\boldsymbol{x}\prec\boldsymbol{x_{ref}}][1-\prod_{\boldsymbol{x}^{*}\in\Omega}\mathbb{1}[\boldsymbol{x}^{*}\nprec\boldsymbol{x}]]\mathrm{d}\boldsymbol{x}.$
Where $\mathbb{1}(\textperiodcentered)$ is the indicator function, which outputs one if its argument is true and zero otherwise, specifically, if $x$ is the newly discovered design that is not dominated by other design points, then the PHV increases. An example is illustrated in Fig.2.
7. Experiment Results
7.1. Overall Analysis
@@
7.1.1. Comparison on Pareto Optimal Set Exploration
Figure 8. Comparison of PHV from IPC-Power optimization with SOTA DSE frameworks.
Figure 9. Comparisons between the Pareto designs in (a) $Perf^{2}/(Power\times Area)$, (b) performance, (c) area, and (d) power.
@@ 7.1.1. Comparison on Pareto Optimal Set Exploration
Table 2. Detailed information of Pareto optimal set search.
@@ Methods
BOOMExplorer 36.94 64 47 (9.4$\times$)
MoDSE 37.04 87 16 (3.2$\times$)
ArchExplorer 38.57 \ 828 (165.6$\times$)
AttentionDSE 39.41 15 5 (1$\times$)
@@
Table 2 shows the exploration results on the 600.perlbench_s dataset across the SOTA DSE frameworks. The “Time” represents the total exploration time of the DSE task. All DSE frameworks, except ArchExplorer, perform a hundred iterations for the Pareto optimal set search. The “Up to 99% PHV Iteration” metric indicates the convergence time of the search, where lower values are better. Experimental results demonstrate that AttentionDSE is 166 $\times$ faster than ArchExplorer, owing to ArchExplorer’s time-consuming instruction analysis method. Compared to regression-based models, AttentionDSE completes the exploration roughly 10$\times$ faster than BOOMExplorer. Even with MoDSE’s use of random sampling, AttentionDSE achieves exploration speeds approximately 3$\times$ faster. Additionally, AttentionDSE converges in fewer iterations, reducing iterations by 75% compared to BOOMExplorer and 83% compared to MoDSE, highlighting the efficiency of the ABA algorithm.
@@
Figure 10. Comparison of MAPE with SOTA DSE frameworks. (a) MAPE of IPC prediction; (b) MAPE of power prediction; (c) MAPE of area prediction; (d) MAPE and normalized inference time with the increased number of parameters on the 600.perlbench_s dataset; (e) The training time changes with the increase in training set size on the 600.perlbench_s dataset.
7.1.2. Comparison on Best Balanced Designs

### v2

@@ V Design of AttentionDSE
In this section, we introduce AttentionDSE, an attention-based framework designed for high-dimensional DSE.
V-A Overview of AttentionDSE
@@
Fig. 3 provides an overview of AttentionDSE. It comprises three primary components: (1) Predictors, which provide rapid and accurate predictions of performance metrics (e.g., IPC) for given micro-architectures across benchmarks; (2) Micro-architecture Serialization, designed to convert various design parameters into input vectors for the predictor while ensuring the positional certainty of parameters; it also reduces the computational burden associated with predictions, thereby expediting the performance evaluation process; (3) Bottleneck Analysis and Iterator, which identify the micro-architectural parameters that introduce performance bottlenecks within micro-architecture modules, facilitating the exploration of promising micro-architectural candidates by selectively modifying these parameters. The process of the AttentionDSE consists of a predictor training stage (①, ②, ③, ④) and an architecture exploration stage (❶, ❷, ❸, ❹, ❺).
Figure 3: Overview of the AttentionDSE.
Figure 4: The workflow of AttentionDSE framework in the exploration stage.
@@
The training stage is a one-time process, and the sequence of operations is represented by hollow numbers. This stage focuses on training predictors for performance metrics, ensuring that the models accurately capture the relationships between micro-architectural parameters and performance metrics. Unlike other approaches that rely on specialized sampling methods, AttentionDSE adopts a random sampling approach and then converts the design point to the sequence input, generating the training set (①, ②, ③). Next, the deep learning model, integrated with the PDA, functioning as the predictor, is trained to predict performance metrics (④). Benefiting from probabilistic attribute correlation in attention calculation, it more easily captures the intricate relationship between parameters, informing hints for the exploration stage.
@@
V-B Attention-based Prediction Model
A fast and accurate performance predictor is essential to replacing traditional simulation-based methods, forming the foundation of an efficient DSE framework. Such predictors enable rapid estimation of performance metrics, significantly accelerating design iterations and reducing overall exploration time. As discussed in Section IV, attention-based neural networks are particularly well-suited for this task: they not only excel at modeling complex, context-dependent relationships in high-dimensional design spaces, but also naturally support parallel computation, making them highly compatible with GPU-based acceleration. Thus, we introduce an attention-based model tailored for single-objective performance prediction. By combining strong modeling capacity with scalable inference, the proposed surrogate supports fast and accurate exploration across large micro-architectural design spaces.
Input: $D$: Design space; $n$: The size of the training set; $k$: Embedding length; DepthDepth: The depth of the model; EpochEpoch: The total epochs for training.
@@ V-B Attention-based Prediction Model
3 Instantiate the prediction token $T$ with embedding length $k$;
@@ 5 Simulate $T$ to obtain $S_{ground\_truth}$ of IPC, Power, and Area;
6 for $epoch\leftarrow 0$ to EpochEpoch do
7 $S_{embed}\leftarrow Embedding(T)$;
8 $S_{embed}\leftarrow Add(S_{embed},PE)$;
9 for $depth\leftarrow 0$ to DepthDepth do
@@ 11    $S_{mlp}\leftarrow MLP(S_{attn})$;
14 $P_{t},Attn\_{score}\leftarrow SelfAttention(S_{mlp})$;
15   Extract analysis token $A_{t}\leftarrow Reduce\_{sum}(Attn\_{score})$;
16 $S_{pred}\leftarrow FullyConnectedLayer(P_{t})$;
@@ 17 $BackwardPropagation(S_{pred},S_{ground\_truth})$;
Algorithm 1 outlines the training procedure of our attention-based predictor. The process begins by randomly sampling $n$ design points from the design space $D$ and simulating each to obtain ground truth metrics (IPC, Power, and Area). This procedure comprises the following steps.
  • $\bullet$
    Embedding: Each discrete, heterogeneous microarchitectural parameter is mapped to a vector representation, enabling the model to reason over diverse parameter types in a unified latent space.
  • $\bullet$
    Position Embedding (PE): Positional encodings are added to retain the identity and ordering of each parameter, allowing the attention mechanism to associate learned importance scores with specific microarchitectural features.
  • $\bullet$
    Prediction Token ($P_{t}$): Prediction token aggregates global information via self-attention, and is processed by an MLP to produce final PPA predictions.
  • $\bullet$
    Analysis Token ($A_{t}$): Derived from attention scores from the final self-attention layer, $A_{t}$ captures the relative importance of each parameter, enabling bottleneck analysis and interpretability.
By embedding structural context and aligning parameter positions, the model effectively captures complex interdependencies among architectural parameters—addressing the key challenge of scalability and insight in high-dimensional DSE.
V-C Perception-driven Attention Mechanism
@@
In previous work [45, 42, 13, 3], the order of parameters is not a concern. Because GPR [45, 3] relies on the distance and similarity between input data rather than their arrangement order in statistical regression methods. Similarly, in ensemble learning [13, 42], the training process of multiple weak learners introduces randomness, making the order of parameters irrelevant. To address this challenge in the attention-based predictor, we introduce a serialization approach termed Perception-driven Serialization (PDS), which preserves architectural information and establishes a foundation for the sliding window attention technique. This method leverages a graph-based framework to map relationships among parameters within each pipeline stage, rooted in a key insight: CPU instruction and data flows are processed sequentially across pipeline stages, with parameters within the same stage interacting far more frequently than those across different stages.
Figure 5: The Perception-driven Serialization approach. (a) A Sample OoO CPU core; (b) Rename Stage; (c) Stage abstraction; (d) The steps of Perception-driven Serialization approach.
An example is illustrated in Fig. 5(a). We chose the Rename stage, shown in Fig. 5(b), as an example, where there are four parameters to be serialized, including the free list (F), the rename map (R), the scoreboard (S), and the history buffer (H). The PDS first transfers the micro-architecture into a perceptual graph, which is an undirected graph based on the dataflow among these components, as depicted in Fig. 5(c). In a perceptual graph, vertices represent the parameters within each pipeline stage, while edges denote the datapaths between these components. The edges are categorized into two types: red edges, which connect to other pipeline stages and are labeled as external edges, and black edges, which connect to components within the current pipeline stage and are labeled as internal edges. To quantify the perception of each component, we leverage the perception degree $D$ of each parameter in the perceptual graph. $D$ is calculated as follows:
$D(\text{parameter})=\sum_{e\in E_{\text{parameter}}^{\text{internal}}}e-\sum_{e\in E_{\text{parameter}}^{\text{external}}}e$ (4)
@@
Figure 6: The attention weight calculation. (a) Vanilla self-attention mechanism; (b) PDA mechanism.
In the clustered sequence, the correlation between adjacent parameters is significantly higher than that between non-adjacent parameters. At this point, the sliding window only needs to cover a local area to capture most of the key dependencies, without the need to calculate global attention. Fig. 6 shows an example of attention weight calculation compared with vanilla self-attention and PDA. For vanilla self-attention, every parameter must calculate the attention weight with every other parameter, as shown in Fig. 6(a). To optimize this process, we set the window size to the maximum perception degree among all parameters. In this example, the window size is set to five, as illustrated in Fig. 6(b).
In general, with the help of the PDA, each parameter calculates the attention weight only with its relative parameters. In this way, for a parameter vector of length $n$, the computational complexity is reduced from $O(n^{2})$ to $O(n)$. Thus, the PDA approach alleviates the limitations posed by the number of parameters, enhances scalability, and reduces the training and inference time of the prediction model.
V-D Attention-aware Bottleneck Analysis
Bottleneck analysis is a crucial step in CPU design, enabling architects to identify the key microarchitectural factors that significantly influence performance. To support this capability, we propose the ABA algorithm (Algorithm 2), which captures intricate relationships between microarchitectural parameters and performance metrics. By identifying the parameters that contribute most to performance variation, ABA effectively guides DSE by prioritizing high-impact parameters, thereby improving both exploration efficiency and design quality.
To enable such analysis in an interpretable and quantifiable manner, we extract a dedicated analysis token $A_{t}$ from the previously introduced prediction model. This token is designed to assess the importance of each input parameter with respect to performance outcomes. A central requirement for ensuring the reliability of this mechanism is maintaining a strict one-to-one correspondence between attention weights and the input parameters. To ensure this, the model preserves token position consistency across all components: an embedding layer, PDA layer, MLP, self-attention, and fully connected (FC) layer. Specifically, the embedding layer applies independent projections to each parameter token, the PDA layer adds positional encoding while preserving order, and both the MLP and self-attention layers operate token-wise to retain the direct mapping between tokens and their associated attention weights.
This rigorous preservation of token alignment is vital: it enables each attention weight to be directly and explicitly mapped back to a specific design parameter. Leveraging this property, the ABA algorithm interprets the attention-weight heatmap generated by the model to identify performance-critical parameters. In this heatmap, each row indicates how much a particular parameter is influenced by others, while each column—indexed by the analysis token $A_{t}$—reflects the degree to which a parameter affects all other parameters. The column-wise sum of attention weights thus serves as a proxy for a parameter’s global impact on performance prediction.
Based on this insight, ABA detects bottleneck parameters by analyzing $A_{t}$: a low value for IPC (indicating high sensitivity) or a high value for Power or Area (indicating dominant cost contribution) signifies that the corresponding parameter is a bottleneck for that specific objective. ABA then updates the design configuration accordingly, refining the Pareto frontier and facilitating iterative improvements in both design quality and exploration efficiency.
@@ 5   Set search queue $Q\leftarrow\varnothing$;
7    $A_{t}\leftarrow Predictor(x)$;
8    $x^{\prime}\leftarrow BottleneckAnalysis(x,A_{t})$;
@@ 16 $\Omega\leftarrow GetParetoOptimal(Predictor(Q))$;
Figure 7: The example of ABA algorithm updating the parameters.
@@ VI Experimental Setup
Simulation. We utilize GEM5 [4] as the timing-accurate simulator and McPAT [23] as the power and area modeling tool. McPAT is configured with technology parameters aligned with industry-standard VLSI design rules, enabling it to estimate power (dynamic and leakage) and area based on microarchitectural specifications, which is consistent with prior DSE works [42, 3]. We extend GEM5 using Python to generate different CPU cores with varying configurations. To bolster the credibility of our experiments, we employ SPEC CPU2017 [37] as the evaluation benchmark. For identifying resource utilization status, we utilize Simpoints [17] for each workload evaluation. Each workload is divided into at most 30 clusters, with each cluster containing ten million instructions.
Baselines. Our baselines include the ActBoost [22], the BOOMExplorer [3], and the MoDSE [42]. Additionally, we implement ArchExplorer [2] for comparison with our ABA algorithm. The baselines represent recent SOTA work. All experiments are conducted on the Out-of-Order (OoO) CPU implemented by Gem5. The design space of the OoO processor is listed in Table I. The values in the third column are formatted as “start number: end number: stride”. A two-level cache hierarchy with an 8192MB DRAM main memory configuration is employed. The size of the design space exceeds $6.89\times 10^{35}$, making exhaustive search impractical.
TABLE I: Micro-architecture design space specification.
@@
 $\text{ MAPE }=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{\left|y_{i}-y_{i}^{\text{real }}\right|}{y_{i}^{\text{real }}}\right)\times 100.$ (5)
@@
$\mathrm{PHV}(\Omega)=\int_{\mathrm{X}}\mathbb{1}[\boldsymbol{x}\prec\boldsymbol{x_{ref}}][1-\prod_{\boldsymbol{x}^{*}\in\Omega}\mathbb{1}[\boldsymbol{x}^{*}\nprec\boldsymbol{x}]]\mathrm{d}\boldsymbol{x}.$ (6)
Where $\mathbb{1}(\textperiodcentered)$ is the indicator function, which outputs one if its argument is true and zero otherwise, specifically, if $x$ is the newly discovered design that is not dominated by other design points, then the PHV increases. See Fig. 2 for an example.
VII Experiment Results
VII-A Overall Analysis
@@
VII-A1 Comparison on Pareto Optimal Set Exploration
Figure 8: Comparison of PHV from IPC-Power optimization with SOTAs.
Figure 9: Comparisons between the Pareto designs in (a) $Perf^{2}/(Power\times Area)$, (b) performance, (c) area, and (d) power.
@@ VII-A1 Comparison on Pareto Optimal Set Exploration
TABLE II: Detailed information of Pareto optimal set search.
@@ Methods
[…]
@@
[…]
@@
Figure 10: Comparison of MAPE with SOTA DSE frameworks. (a) MAPE of IPC prediction; (b) MAPE of power prediction; (c) MAPE of area prediction; (d) MAPE and normalized inference time with the increased number of parameters on the 600.perlbench_s dataset; (e) The training time changes with the increase in training set size on the 600.perlbench_s dataset.
VII-A2 Comparison on Best Balanced Designs